Show that the function f defined by f(x)= 1/x^2 is uniformly continuous on [a, ∞] , a> 0

1
Expert's answer
2022-02-09T17:32:28-0500

For a function "f: A \rightarrow B" to be uniformly continuous we need that

"\forall \varepsilon>0 \exists \delta>0 \forall x, y \in A:|x-y|<\delta \Rightarrow|f(x)-f(y)|<\varepsilon"

We have in your case "f(x)=\frac{1}{x^{2}}."

We have "\forall x, y \in[1, \infty)"

"|f(x)-f(y)|=\left|\frac{1}{x^{2}}-\frac{1}{y^{2}}\right|=\left|\frac{y^{2}-x^{2}}{x^{2} y^{2}}\right|=\frac{1}{x^{2} y^{2}}\left|y^{2}-x^{2}\right|=\frac{1}{x^{2} y^{2}}|y+x||y-x|"

Now, since "y+x>0" and "|y-x|=|x-y|" we get

"\frac{1}{x^{2} y^{2}}|y+x \| y-x|=\frac{x+y}{x^{2} y^{2}}|x-y|=\left(\frac{1}{x y^{2}}+\frac{1}{x^{2} y}\right)|x-y|<|x-y|"

since "\frac{1}{x y^{2}}<\frac{1}{2}" and "\frac{1}{x^{2} y}<\frac{1}{2}\ \text{for}\ x, y \in[1, \infty)."

Therefore if you set "\delta=\varepsilon" we get

"|f(x)-f(y)|<|x-y|<\delta=\varepsilon"